# German Research Center for Artificial Intelligence, Kaiserslautern, Germany

### v1

@@ 5.0.1 Acknowledgements
Appendix 0.A Experimental setting
Regarding the implementation details and preprocessing, we apply a z-score normalization to the input data and encode the categorical and ordinal features (like land-cover and canopy height) with a one-hot-vector. We use a 1D CNN for the temporal features, an MLP encoder for the static features and as a prediction head. We use two layers on all architectures with 128 units and a 20% of dropout. We use the ADAM optimizer with a batch-size of 128, early stopping, and standard loss functions (cross-entropy in classification and mean squared error in regression).
In Table 3 we display the number of features per sensor in each dataset.
@@ Appendix 0.B Predictive performance results
As a complement to the PRS results shown in the main content of the manuscript, we include predictive performance scores, the F1 in classification and $R^{2}$ in regression. Fig. 7, 8, 9 show the results for the CH, LFMC, and PM25 datasets respectively. In classification, the ensemble-based methods are the best in the F1 score. However, in regression, the input-level fusion methods are the best in the $R^{2}$ score, while for missing sensor cases, the proposed methods with SensD are the best options.
@@
Appendix 0.C Extended ablation results
In Fig 10 and 11 the predictive performance ($R^{2}$) is shown at different levels of the SensD ratio value in the LFMC and PM25 datasets.
Figure 10: $R^{2}$ scores of the ISensD method with different SensD ratio values in the LFMC data. The dashed lines show the no ratio version of the ISensD method.
Figure 11: $R^{2}$ scores of the ISensD method with different SensD ratio values in the PM25 data. The dashed lines show the no ratio version of the ISensD method.
In Table 4 and 5 the predictive performance ($R^{2}$) is shown for different configurations of the ESensI model in the LFMC and PM25 datasets.
Table 4: $R^{2}$ results of different configurations in the ESensI model for the PM25 data. The results are obtained in the full-sensor evaluation.
Shared weights Sensor encoding Normalization Others $R^{2}$ score
- - - $0.313$
$\checkmark$ - - $0.281$
$\checkmark$ $\checkmark$ - $0.309$
$\checkmark$ $\checkmark$ $\checkmark$ addition $0.317$
$\checkmark$ $\checkmark$ $\checkmark$ concatenation $0.297$
Table 5: $R^{2}$ results of different configurations in the ESensI model for the LFMC data. The results are obtained in the full-sensor evaluation.
Shared weights Sensor encoding Normalization Others $R^{2}$ score
- - - $0.308$
$\checkmark$ - - $0.281$
$\checkmark$ $\checkmark$ - $0.274$
$\checkmark$ $\checkmark$ $\checkmark$ addition $0.241$
$\checkmark$ $\checkmark$ $\checkmark$ concatenation $0.196$

### v2

@@ 6.0.1 Acknowledgements
Appendix 0.A Dataset setting
In Table 3 we display the number of features per sensor in each dataset. Regarding the data preprocessing, we apply a z-score normalization to the input data and encode the categorical and ordinal features (like land-cover and canopy height) with a one-hot-vector.
@@ Appendix 0.B Predictive performance results
As a complement to the PRS results shown in the main content of the manuscript, we include predictive performance scores, the F1 in classification, and $R^{2}$ in regression. Fig. 7, 8, 9 show the results for the CH, LFMC, and PM25 datasets respectively. In classification, the ensemble-based methods are the best in the F1 score. However, in regression, the input-level fusion methods are the best in the $R^{2}$ score, while for missing sensor cases, the proposed methods with SensD are the best options.
@@
Appendix 0.C Alternative encoders
Figure 10: Predictive performance with different encoders in the CH data.
Figure 11: Predictive performance with different encoders in the PM25 data.
Figure 12: Predictive performance with different encoders in the LFMC data.
In Fig. 10, 11, and 12,ex we present the predictive performance results when the encoder architecture in the proposed models is changed. The options explored are TempCNN (the default we use in our paper, 1D CNN applied over time), Temporal Attention Encoder (TAE), and Gated Recurrent Unit (GRU)-based RNN. In some cases, the TempCNN architecture is more robust, like in the CH data. While in others, the GRU architecture is more robust, like in the PM25 data. However, we can see that the main pattern is the same for all encoder architectures, with no extreme changes in the overall results.